OpenStudy (anonymous):

An Ice Cube machine makes 24 cubes every 1/2 hour. At This rate, how many ice cubes can it make in 45 minutes?

OpenStudy (turingtest):

#cubes=(cubes/hour)(#hours) C=rt 24=r(1/2) r=48 cubes/hr in 45 min=3/4hrs C=rt=48(3/4)=36 cubes

OpenStudy (anonymous):

So the answer is 36 cubes?

OpenStudy (turingtest):

yes